what is the sine of angle A
the triangle is a right angle triangle with a hypotenuse of 10 , opposite 6, and adjacent 8
angle A is between 10 and 8...

Answer by chibisan(131)

draw out a triangle to make it more clearer.
(to make a spacing , i drew the dots :} )
....l\
8 l \10
....l__\
.......6

sinA = opp/hyp = 6/10 = 3/5